# ProtChatGPT: Towards Understanding Proteins with Large Language Models

### v1

@@ Appendix A Details on Protein-Language Representation Learning
Obtaining queries that can extract informative protein representation regarding text is significant for protein-language alignment. To achieve this, we connect our PLP-former with the ESM-1b (Rives et al., 2021) model during the representation learning phase and train with the protein-language pairs. Following (Li et al., 2022a, 2023), we jointly train our model with three distinct pre-training tasks: Protein-Text Contrastive learning (PTC), Protein-grounded Text Generation (PTG), and Protein-Text Matching (PTM). Although these tasks utilize the same model structure and input format, they differ in the attention masking strategy applied between queries and text, thereby modulating their interaction.
Protein-Text Contrastive Learning (PTC).
@@
For Protein-Text Contrastive Learning, by maximizing the mutual information, we aim to ensure the latent representation of protein and text are well-aligned. Specifically, given the query representation from the protein transformer $\boldsymbol{E}_{seq}$, we align it with the corresponding text embedding $t$. This is achieved by maximizing the similarity of positive pairs against those negative pairs where we directly use the embedding of $[cls]$ tokens from the text transformer as $t$. Given that the output of the protein transformer comprises multiple embeddings, we calculate the pairwise similarity between each query output and $t$. We then choose the highest value to represent the protein-text similarity. To prevent any information leakage, we utilize an unimodal self-attention mask that restricts direct interaction between queries and text.
Protein-grounded Text Generation (PTG).
@@
The PTG task is designed to ensure that the learned queries can efficiently derive text-relevant information from the protein sequence. To accomplish this, we train the PLP-former to produce descriptions matching the respective protein sequences. Since the PLP-former prevents direct interaction between the frozen ESM-1b and text tokens, the data needed for description generation must first be garnered by the queries, ensuring efficient information extraction. We use a multimodal causal self-attention mask, to manage the interaction between queries and text. While queries can interact with one another, they cannot engage with the text tokens. Conversely, each text token can reference all queries as well as its preceding text tokens. Additionally, we substitute the $[CLS]$ token with a $[DEC]$ token at the beginning of the text sequence to indicate the decoding task.
Protein-Text Matching (PTM).
@@
Protein-Text Matching task is leveraged for fine-grained protein-text representation alignment. This task is designed as a binary classification task where the model needs to determine if a given image-text pair aligns (positive) or misaligns (negative). We employ a bi-directional self-attention mask, allowing all queries and texts to mutually attend. As a result, the obtained query embeddings, $\boldsymbol{E}_{seq}$, encompass multimodal information. Each of these embeddings is then passed through a binary linear classifier to derive a logit, with the final matching score being the average of logits across all queries. For crafting informative negative pairs, we utilize the hard negative mining technique as described in (Li et al., 2021).
Appendix B Details and Examples of the Training Datasets
@@
ProtDescribe Dataset. ProtDescribe dataset (Xu et al., 2023) is designed to augment protein sequences with text descriptions of their functions and other important properties. This dataset contains 549,000 pairs of proteins, including the EntryName, ProteinName, Function (obtained from Uniport), SubcellularLocation, Similarity, and ProteinSequence. In our implementation, we only use the protein sequence and the corresponding protein functions for the training of PLP-former. The PLP-former takes both sequence and descriptions as inputs, after three combined training tasks (Appendix. A, then outputs 32 learned tokens indicating the sequence embeddings that are most relevant to the descriptions. Some examples of this dataset are given in Table. 3.
Table 2: Examples of RCSB-PDB dataset with the protein entry id from Protein Data Bank.
@@
Has immunoglobulin-binding and
hemagglutination properties,
and can bind to mannose.
Essential for virulence. May be
involved in LPS biosynthesis
or polysaccharide transport.
@@
Plays a role in virus cell tropism,
and may be required for efficient
virus replication in macrophages.
@@
Belongs to the asfivirus MGF 110 family.
@@
Table 3: Examples of ProtDescribe dataset.
@@
The x-ray crystal structure of dimeric (+)-bornyl diphosphate synthase, a metal-requiring monoterpene cyclase from Salvia officinalis, is reported at
2.0-A resolution. Each monomer contains two alpha-helical domains: the C-terminal domain catalyzes the cyclization of geranyl diphosphate,
orienting and stabilizing multiple reactive carbocation intermediates; the N-terminal domain has no clearly defined function, although its N
terminus caps the active site in the C-terminal domain during catalysis. Structures of complexes with aza analogs of substrate and carbocation
intermediates, as well as complexes with pyrophosphate and bornyl diphosphate, provide "snapshots" of the terpene cyclization cascade.
@@
The role of hydrophobic amino acids in the formation of hydrophobic cores as one of the major driving forces in protein folding has been extensively
studied. However, the implication of neutral solvent-exposed amino acids is less clear and available information is scarce. We have used a combinatorial
approach to study the structural relevance of three solvent-exposed residues (Tyr(327), Thr(329), and Gln(331)) located in thebeta-sheet of the
tetramerization domain of the tumor suppressor p53 (p53TD). A conformationally defined peptide library was designed where these three positions
were randomized. The library was screened for tetramer stability. A set of p53TD mutants containing putative stabilizing or destabilizing residue
combinations was synthesized for a thermodynamic characterization. Unfolding experiments showed a wide range of stabilities, with T(m) values
between 27 and 83 degrees C. Wild type p53TD and some highly destabilized and stabilized mutants were further characterized. Thermodynamic
and biophysical data indicated that these proteins were folded tetramers, with the same overall structure, in equilibrium with unfolded monomers.
An NMR study confirmed that the main structural features of p53TD are conserved in all the mutants analyzed. The thermodynamic stability of
the different p53TD mutants showed a strong correlation with parameters that favor formation and stabilization of the beta-sheet. We propose
that stabilization through hydrophobic interactions of key secondary structure elements might be the underlying mechanism for the strong influence
of solvent-exposed residues in the stability of p53TD.
@@
Two crystal modifications of a collagen model peptide, (Pro-Pro-Gly)(4)-Hyp-Hyp-Gly-(Pro-Pro-Gly)(4) [where Hyp is (4R,2S)-L-hydroxyproline],
@@
showed very similar unit-cell parameters and belonged to the same space group P2(1). Both crystals exhibited pseudo-merohedral twinning. The
@@
main difference was in their molecular-packing arrangements. One modification showed pseudo-hexagonal packing, while the other showed
@@
pseudo-tetragonal packing. Despite their different packing arrangements, no significant differences were observed in the hydration states of these
@@
modifications. The peptide in the pseudo-tetragonal crystal showed a cyclic fluctuation of helical twists with a period of 20 A, while that in the
@@
pseudo-hexagonal crystal did not. In these modifications, the puckering conformations of four of the 12 Hyp residues at the X position of the
@@
Hyp(X)-Hyp(Y)-Gly sequence were in the opposite conformations to the previous hypothesis that Hyp(X) residues involved in Hyp(X):Hyp(Y)
and Hyp(X):Pro(Y) stacking pairs prefer up-puckering and down-puckering conformations, respectively. Detailed investigation of the molecular
@@
interactions between Hyp(X) and adjacent molecules revealed that these opposite conformations appeared because the puckering conformation,
which follows the hypothesis, is subject to steric hindrance from the adjacent molecule.
@@
RCSB-PDB Dataset. RCSB-PDB dataset (Guo et al., 2023) is originally sourced from the Research Collaboratory for Structural Bioinformatics Protein Data Bank  [^1] https://www.rcsb.org, which includes 204,826 experimentally determined 3D structures. ProteinChat (Guo et al., 2023) further selected 143,508 proteins that have a primary publication linked with a PubMed ID and can be extracted with a valid chain. In our implementation, we compile both 1D amino acid sequences and 3D atomic coordinates, along with the abstract of corresponding scientific literature for the second-stage training. With the PLP-former and structure encoder frozen, LLM takes out the sequence embedding, structure embedding and the user question as inputs, and produces the related answers. Some examples of this dataset are given in Table. 2.
Appendix C Details on Metrics

### v2

@@ Appendix E Details on Protein-Language Representation Learning
Obtaining queries that can extract informative protein representation regarding text is significant for protein-language alignment. To achieve this, we connect our PLP-former with the ESM-1b (Rives et al., 2021) model during the representation learning phase and train with the protein-language pairs. Following (Li et al., 2022a, 2023), we jointly train our model with three distinct pretraining tasks: Protein-Text Contrastive learning (PTC), Protein-grounded Text Generation (PTG), and Protein-Text Matching (PTM). Although these tasks utilize the same model structure and input format, they differ in the attention masking strategy applied between queries and text, thereby modulating their interaction.
Protein-Text Contrastive Learning (PTC)
@@
For Protein-Text Contrastive Learning, by maximizing the mutual information, we aim to ensure the latent representation of protein and text are well-aligned. Specifically, given the query representation from the protein transformer ${\bm{E}}_{seq}$, we align it with the corresponding text embedding $t$. This is achieved by maximizing the similarity of positive pairs against those negative pairs where we directly use the embedding of $[cls]$ tokens from the text transformer as $t$. Given that the output of the protein transformer comprises multiple embeddings, we calculate the pairwise similarity between each query output and $t$. We then choose the highest value to represent the protein-text similarity. To prevent any information leakage, we utilize an unimodal self-attention mask that restricts direct interaction between queries and text.
Protein-grounded Text Generation (PTG)
@@
The PTG task is designed to ensure that the learned queries can efficiently derive text-relevant information from the protein sequence. We train the PLP-former to produce descriptions matching the respective protein sequences. Since the PLP-former prevents direct interaction between the frozen ESM-1b and text tokens, the data needed for description generation must first be garnered by the queries, ensuring efficient information extraction. We use a multimodal causal self-attention mask, to manage the interaction between queries and text. While queries can interact with one another, they cannot engage with the text tokens. Conversely, each text token can reference all queries as well as its preceding text tokens. Additionally, we substitute the $[CLS]$ token with a $[DEC]$ token at the beginning of the text sequence to indicate the decoding task.
Protein-Text Matching (PTM)
@@
The protein-text Matching task is leveraged to align fine-grained protein-text representations. This task is designed as a binary classification task where the model needs to determine if a given image-text pair aligns (positive) or misaligns (negative). We employ a bi-directional self-attention mask, allowing all queries and texts to attend mutually. As a result, the obtained query embeddings, ${\bm{E}}_{seq}$, encompass multimodal information. Each of these embeddings is then passed through a binary linear classifier to derive a logit, with the final matching score being the average of logits across all queries. For crafting informative negative pairs, we utilize the hard negative mining technique as described in (Li et al., 2021).
Appendix F Details and Examples of the Training Datasets
@@
ProtDescribe Dataset. The ProtDescribe dataset (Xu et al., 2023) is designed to augment protein sequences with text descriptions of their functions and other important properties. This dataset contains 549,000 pairs of proteins, including the EntryName, ProteinName, Function (obtained from Uniport), SubcellularLocation, Similarity, and ProteinSequence. In our implementation, we only use the protein sequence and the corresponding protein functions for the training of PLP-former. The PLP-former takes both sequence and descriptions as inputs, after three combined training tasks (Appendix. E), then outputs 32 learned tokens indicating the sequence embeddings that are most relevant to the descriptions. Some examples of this dataset are given in Table. 4.
Table 4. Examples of the original ProtDescribe dataset.
@@
Has immunoglobulin-binding
and hemagglutination
properties, and can bind to
mannose. Essential for
virulence. May be involved in
LPS biosynthesis or
polysaccharide transport.
@@
Plays a role in virus cell
tropism, and may be required
for efficient virus replication
in macrophages.
@@
Belongs to the
asfivirus MGF
110 family.
@@
Table 5. Examples of the RCSB-PDB dataset with the protein entry ID from Protein Data Bank.
@@
The x-ray crystal structure of dimeric (+)-bornyl diphosphate synthase, a metal-requiring monoterpene cyclase from Salvia
officinalis, is reported at 2.0-A resolution. Each monomer contains two alpha-helical domains: the C-terminal domain
catalyzes the cyclization of geranyl diphosphate, orienting and stabilizing multiple reactive carbocation intermediates;
the N-terminal domain has no clearly defined function, although its N terminus caps the active site in the C-terminal domain
during catalysis. Structures of complexes with aza analogs of substrate and carbocation intermediates, as well as complexes
with pyrophosphate and bornyl diphosphate, provide ”snapshots” of the terpene cyclization cascade.
@@
The role of hydrophobic amino acids in the formation of hydrophobic cores as one of the major driving forces in protein
folding has been extensively studied. However, the implication of neutral solvent-exposed amino acids is less clear and
available information is scarce. We have used a combinatorial approach to study the structural relevance of three solvent
-exposed residues (Tyr(327), Thr(329), and Gln(331)) located in thebeta-sheet of the tetramerization domain of the tumor
suppressor p53 (p53TD). A conformationally defined peptide library was designed where these three positions were
randomized. The library was screened for tetramer stability. A set of p53TD mutants containing putative stabilizing or
destabilizing residue combinations was synthesized for a thermodynamic characterization. Unfolding experiments showed
a wide range of stabilities, with T(m) values between 27 and 83 degrees C. Wild type p53TD and some highly destabilized
and stabilized mutants were further characterized. Thermodynamic and biophysical data indicated that these proteins were
folded tetramers, with the same overall structure, in equilibrium with unfolded monomers. An NMR study confirmed that
the main structural features of p53TD are conserved in all the mutants analyzed. The thermodynamic stability of the different
p53TD mutants showed a strong correlation with parameters that favor the formation and stabilization of the beta-sheet. We
propose that stabilization through hydrophobic interactions of key secondary structure elements might be the
underlying mechanism for the strong influence of solvent-exposed residues in the stability of p53TD.
@@
Two crystal modifications of a collagen model peptide, (Pro-Pro-Gly)(4)-Hyp-Hyp-Gly-(Pro-Pro-Gly)(4) [where Hyp is
@@
(4R,2S)-L-hydroxyproline], showed very similar unit-cell parameters and belonged to the same space group P2(1). Both
@@
crystals exhibited pseudo-merohedral twinning. The main difference was in their molecular-packing arrangements. One
modification showed pseudo-hexagonal packing, while the other showed pseudo-tetragonal packing. Despite their
@@
different packing arrangements, no significant differences were observed in the hydration states of these modifications.
@@
The peptide in the pseudo-tetragonal crystal showed a cyclic fluctuation of helical twists with a period of 20 A, while that
@@
in the pseudo-hexagonal crystal did not. In these modifications, the puckering conformations of four of the 12 Hyp residues
@@
at the X position of the Hyp(X)-Hyp(Y)-Gly sequence were in the opposite conformations to the previous hypothesis that
Hyp(X) residues involved in Hyp(X):Hyp(Y) and Hyp(X):Pro(Y) stacking pairs prefer up-puckering and down-puckering
conformations, respectively. Detailed investigation of the molecular interactions between Hyp(X) and adjacent molecules
@@
revealed that these opposite conformations appeared because the puckering conformation, which follows the hypothesis,
is subject to steric hindrance from the adjacent molecule.
Table 6. An example of our instruction tuning dataset. We only show the input abstract and the generated instruction pairs for simplicity.
Abstract Instruction pairs
Assembly of polymerase chain reactions
at room temperature can sometimes lead
to low yields or unintentional products
due to mispriming. Mutation of isoleucine
707 to leucine in DNA polymerase I from
Thermus aquaticus substantially decreases
its activity at room temperature without
compromising its ability to amplify DNA.
To understand why a conservative change to
the enzyme over 20 u00c5 from the active
site can have a large impact on its activity
at low temperature, we solved the X-ray
crystal structure of the large (5’-to-3’
exonuclease-deleted) fragment of Taq DNA
polymerase containing the cold-sensitive
mutation in the ternary (E-DNA-ddNTP) and
binary (E-DNA) complexes. The I707L
KlenTaq1 ternary complex was identical to the
wild-type in the closed conformation except
for the mutation and a rotamer change in
nearby phenylalanine 749, suggesting that the
enzyme should remain active. However, soaking
out of the nucleotide substrate at low temper-
ature results in an altered binary complex made
possible by the rotamer change at F749 near the
tip of the polymerase O-helix. Surprisingly, two
adenosines in the 5’-template overhang fill the
vacated active site by stacking with the primer
strand, thereby blocking the active site at low
temperature. Replacement of the two overhanging
adenosines with pyrimidines substantially
increased activity at room temperature by keeping
the template overhang out of the active site,
confirming the importance of base stacking. These
results explain the cold-sensitive phenotype of the
I707L mutation in KlenTaq1 and serve as an example
of a large conformational change affected by a
conservative mutation.
Q: What is the effect of mutating isoleucine 707 to leucine in DNA polymerase
I from Thermus aquaticus?
A: The mutation substantially decreases the enzyme’s activity at room
temperature without compromising its ability to amplify DNA.
Q: Why is it significant that the mutation site is over 20 Å away from
the active site?
A: It is significant because it demonstrates that even conservative changes
far from the active site can have a large impact on enzyme activity at low
temperatures.
Q: What method was used to understand the impact of the I707L mutation on
enzyme activity?
A: The X-ray crystal structure of the large fragment of Taq DNA polymerase
containing the cold-sensitive mutation was solved in both ternary
(E-DNA-ddNTP) and binary (E-DNA) complexes.
Q: How did the I707L mutation affect the structure of the ternary complex?
A: The I707L KlenTaq1 ternary complex was identical to the wild-type in the
closed conformation, except for the mutation and a rotamer change in nearby
phenylalanine 749.
Q: How does the I707L mutation affect the enzyme’s activity at room
temperature?
A: The I707L mutation substantially decreases the enzyme’s activity at room
temperature.
Q: What happens to the I707L mutation’s nucleotide substrate at low
temperatures?
A: The nucleotide substrate soaks out at low temperatures, resulting in an
altered binary complex.
Q: What specific mutation was made to DNA polymerase I?
A: Isoleucine 707 was mutated to leucine in DNA polymerase I.
Q: What structural method was used to analyze the I707L mutation’s impact?
A: X-ray crystallography was used to determine the structure of the mutated
enzyme.
Q: What was observed in the ternary complex structure of the I707L mutated
enzyme?
A: The ternary complex structure was identical to the wild-type except for the
mutation and a rotamer change in phenylalanine 749.
Q: What effect does the rotamer change at phenylalanine 749 have on the
enzyme?
A: The rotamer change enables two adenosines from the 5’-template overhang
to fill the vacated active site, blocking it at low temperatures.
Table 7. Another example of our instruction tuning dataset.
Abstract Instruction pairs
The bacterial ubiD and ubiX or the
homologous fungal fdc1 and pad1
genes have been implicated in the
non-oxidative reversible decarboxy-
lation of aromatic substrates, and
play a pivotal role in bacterial ubi-
quinone (also known as coenzyme
Q) biosynthesis or microbial bio-
degradation of aromatic compounds,
respectively. Despite biochemical
studies on individual gene products,
the composition and cofactor require-
ment of the enzyme responsible for
in vivo decarboxylase activity
remained unclear. Here we show that
Fdc1 is solely responsible for the
reversible decarboxylase activity, and
that it requires a new type of cofactor:
a prenylated flavin synthesized by the
associated UbiX/Pad1. Atomic resolu-
tion crystal structures reveal that two
distinct isomers of the oxidized co-
factor can be observed, an isoalloxazine
N5-iminium adduct and a N5 secondary
ketimine species with markedly altered
ring structure, both having azomethine
ylide character. Substrate binding posi-
tions the dipolarophile enoic acid group
directly above the azomethine ylide
group. The structure of a covalent
inhibitor-cofactor adduct suggests that
1,3-dipolar cycloaddition chemistry
supports reversible decarboxylation in
these enzymes. Although 1,3-dipolar
cycloaddition is commonly used in
organic chemistry, we propose that this
presents the first example, to our know-
ledge, of an enzymatic 1,3-dipolar
cycloaddition reaction. Our model for
Fdc1/UbiD catalysis offers new routes
in alkene hydrocarbon production or
aryl (de)carboxylation.
Q: What enzymatic activity is Fdc1 responsible for?
A: Fdc1 is responsible for the reversible decarboxylation of aromatic substrates.
Q: What type of cofactor does Fdc1 require for its activity?
A: Fdc1 requires a prenylated flavin cofactor for its activity.
Q: How is the cofactor for Fdc1 synthesized?
A: The cofactor for Fdc1 is synthesized by UbiX/Pad1.
Q: What are the two isomeric forms of the oxidized cofactor in Fdc1?
A: The two isomeric forms are the isoalloxazine N5-iminium adduct and the N5
secondary ketimine species.
Q:What characterizes the cofactor forms in Fdc1?
A: Both cofactor forms in Fdc1 have azomethine ylide character.
Q: How does substrate binding influence Fdc1’s catalytic mechanism?
A: Substrate binding positions the enoic acid group directly above the azomethine
ylide group, facilitating the catalytic process.
Q: What does the covalent inhibitor-cofactor adduct structure in Fdc1 suggest about
its mechanism?
A: It suggests that 1,3-dipolar cycloaddition chemistry is a mechanism supporting
reversible decarboxylation in Fdc1.
Q: What makes Fdc1’s reaction significant in enzymatic chemistry?
A: Fdc1’s reaction may represent the first enzymatic example of 1,3-dipolar cycloaddition,
according to the text.
Q: What potential industrial application does Fdc1’s enzymatic activity suggest?
A: Fdc1’s enzymatic activity suggests potential applications in alkene hydrocarbon
production and aryl (de)carboxylation processes.
Q: What does the altered ring structure of the cofactor in Fdc1 indicate about its function?
A: The altered ring structure indicates significant changes in electronic structure that
likely influence the enzyme’s catalytic activity.
@@
RCSB-PDB Dataset. The RCSB-PDB dataset (Guo et al., 2023) is originally sourced from the Research Collaboratory for Structural Bioinformatics Protein Data Bank  [^1] https://www.rcsb.org, which includes 204,826 experimentally determined 3D structures. ProteinChat (Guo et al., 2023) further selected 143,508 proteins that have a primary publication linked with a PubMed ID and can be extracted with a valid chain. Some examples of this dataset are given in Table. 5.
Instruction Tuning Dataset. Our refined instruction tuning dataset is based on the RCSB-PDB dataset. The RCSB-PDB dataset contains a protein entry ID from Protein Data Bank, a pdb file with atomic coordinates representing the molecular structure, along with the abstract of corresponding scientific literature. In this paper, we extract the amino acid chain as the primary sequence and further use the abstract to generate 10 Q&A pairs for instruction tuning. In summary, for every protein taken into account, we compile its residue sequences, secondary structure label, tertiary atomic coordinates, along with corresponding 10 paired Q&A descriptions. We use the pretrained Bonito (Nayak et al., 2024), an open-source model  [^2] https://huggingface.co/BatsResearch/bonito-v1 for conditional task generation from specialized domains across three task types: yes-no question answering, extractive question answering, and natural language inference. Bonito takes the unannotated text and the task attribute as inputs, and the output consists of the instruction and the response. In this paper, we use Bonito to generate synthetic Q&A pairs based on the corresponding scientific literature (only abstract used for simplicity). In our implementation, the task attribute is set as “ extractive question answering”. Some examples of our refined dataset are given in Table. 6 and Table. 7.
Appendix G More Implementation Details
During training, the PLP-former is initialized with the pre-trained weights of PubMedBERT (Gu et al., 2021), whereas the cross-attention layers and learnable tokens are randomly initialized. Following the same setting as PubMedBERT, we use 32 learnable tokens with a dimension of 768, which is the same as the hidden dimension of the PLP-former. We pre-train the PLP-former on the ProtDescribe dataset for 20K epochs with a batch size of 64. We use the AdamW (Loshchilov and Hutter, 2017) optimizer with $\beta_{1}$ = 0.9, $\beta_{2}$ = 0.98, and a weight decay of 0.05. We use a cosine learning rate decay with a peak learning rate of 1e-4, a minimum learning rate of 8e-5, and a linear warm-up of 5K iterations. Note that we apply PLP only to the sequence embeddings since the reported protein structures are much less than sequences. For example, there are 182K experimentally determined structures in PDB (Berman et al., 2000) while 47M protein sequences in Pfam (Mistry et al., 2021). Thus we only use selected sequence-structure pairs from the RCSB-PDB dataset during the following stage. During contrastive learning of our Protein Context Gating (PCG) module, the temperature parameter $\tau$ is empirically set to 0.8, and $k$ is set to 128 as the number of both positive and negative samples. Finally, we freeze the PLP-former, PCG module and LLM decoder, and train the projection adapter for 1K epochs with a batch size of 128. The minimum learning rate for the second stage is set as 5e-5. We also conduct experiments on fine-tuning LLMs, evaluating both full parameter tuning and LoRA (Hu et al., 2021) tuning on Vicuna-13B. Additionally, we fine-tuned LL A MA3 with LoRA for comparison. To avoid out-of-memory issues, we randomly trim excessively long protein sequences and structures to 3,000 amino acids. All experiments are performed on 2 NVIDIA H-100 (80GB) GPUs. Our model with Vicuna-13B requires 3 days for the multi-level protein-language alignment stage and 2.5 days for the instruction tuning stage (full-parameter fine-tuning), while the LL A MA3-8B model takes approximately 1 day for the instruction tuning stage (LoRA fine-tuning). The pre-trained Vicuna-13B and LL A MA3-8B are obtained through the official repositories [^3] https://github.com/meta-llama/llama3, [^4] https://huggingface.co/lmsys/vicuna-7b-v1.5.
Appendix H Additional Experiments
H.1. More results on LLM fine-tuning
In our implementation, we use the frozen LLM encoder as default since the projection adapter works as an information bottleneck between hybrid protein features and the LLM decoder. We further improve the performance by deploying instruction tuning on the LLMs, as described in Section 4.5. We perform 16-bit LoRA (Hu et al., 2021) on Vicuna-13B and quantized 4-bit QLoRA (Dettmers et al., 2024) on LL A MA3-8B. We also perform full-parameter funding on the Vicuna-13B model. Detailed results are shown in Table. 8.
Table 8. Quantitative Comparisons on fine-tuned LLM decoders. $\uparrow$ indicates that a higher value corresponds to better performance. FP means full parameter tuning.
Method BLEU-1 $\uparrow$ BLEU-4 $\uparrow$ ROUGE-L $\uparrow$ METEOR $\uparrow$ CIDEr $\uparrow$ SPICE $\uparrow$
PubMed
BERTScore $\uparrow$
ProtChatGPT (Vicuna-13B) 0.618 0.399 0.493 0.299 0.640 0.321 0.465
ProtChatGPT (LL
A
MA3-8B)
0.643 0.417 0.515 0.326 0.687 0.346 0.487
ProtChatGPT (Vicuna-13B-LoRA) 0.670 0.435 0.547 0.344 0.695 0.377 0.505
ProtChatGPT (LL
A
MA3-8B-QLoRA)
0.698 0.453 0.556 0.354 0.746 0.391 0.529
ProtChatGPT (Vicuna-13B-FP) 0.735 0.479 0.603 0.379 0.761 0.416 0.557
H.2. Experiments on Protein Retrieval
In this paper, we design a Protein Context Gating (PCG) module to align the protein sequence with secondary structure embedding, and further use contrastive learning to align the tertiary structures. To validate the effectiveness of our multi-level protein alignment, we conducted cross-level protein retrieval experiments. Specifically, we randomly selected sequences from the test set, processed them through the PCG module to obtain ${\bm{E}}_{align}$, and then identified the closest protein by calculating the cosine similarity between ${\bm{E}}_{align}$ and the tertiary structure embedding ${\bm{E}}_{ter}$ of other proteins. The results are shown in Figure 6.
Figure 6. Results on cross-level protein retrieval task.
Appendix I Details on Metrics